OpenStudy (anonymous):

You deposit $100 into a savings plan at the end of each year. The interest rate is 6% compounded annually. Find the value of the annuity after 7 years. Do not round until the final answer. Then, round to the nearest cent

OpenStudy (anonymous):

formula for compound interest rate \[A = P (1+r/n)^{nt}\] A = final amount P = initial investment r = interest rate n = number of times the interest is compounded per year t = number of years for this P = 100, r = 0.06, n = 1, t = 7. solve for A

OpenStudy (anonymous):

100 (1+0.06/1)^1 *7 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got the answer 150.3630 but for some reason it's not on the choices ... what should i do >?

OpenStudy (anonymous):

Those are the choices A $2,506.05 B $697.53 C $839.38 D $236.42

OpenStudy (anonymous):

o.o... that's some big numbers there.. let me see what i did wrong >.<

OpenStudy (anonymous):

the answer is c

OpenStudy (anonymous):

Can you tell me how you got it please ?

OpenStudy (anonymous):

\[FV = C ((1+r)^t-1)/r\] Future value of an ordinary annuity FV = future value C = current cash = 100 r = rate = 0.06 t = time = 7

OpenStudy (anonymous):

Thank you so much for the help I just figured another way to do it , it's a lot of steps but i got the same result thank you so much for the help

OpenStudy (anonymous):

o.o how did you do it? *curious*

OpenStudy (anonymous):

Year 1 : 100 Dollars Year 2 : 100 ( 1+0.06) + 100 Year 3 : 318.36 ( 1+0.06) +100 Year 4 : 437.4616 ( 1+0.06)+100 Year 5 : 563.709296 ( 1+0.06)+100 Year 6 : 697.5318538 (1+0.06)+100 Year 7 : 697.5318538 (1+0.06)+100 = 839.383765